OpenStudy (anonymous):

please solve the equation, and show work: 4(x+1)^2=64

sam (.sam.):

Divide each term in the equation by 4. (4(x+1)^(2))/4=64/4 simplify the left and right-hand side of the equation by simplifying each term. (x+1)^(2)=16 Take the square root of each side of the equation to setup the solution for x. sqrt((x+1)^(2))=+-sqrt(16) Remove the perfect root factor (x+1) under the radical to solve for x. (x+1)=+-sqrt(16) ---------------------------------------------------------------- Pull all perfect square roots out from under the radical. In this case, remove the 4 because it is a perfect square. (x+1)=+-4 ---------------------------------------------------------------- First, substitute in the + portion of the +- to find the first solution. (x+1)=4 x=-1+4 x=3 --------------------------------------------------------------- Next, substitute in the - portion of the +- to find the second solution. (x+1)=-4 x=-1-4 x=-5 -------------------------------------------------------------- The complete solution is the result of both the + and - portions of the solution. x=3,-5